OpenStudy (anonymous):

The slope of a line is -4 and its y-intercept is (0,4). What is the equation of the line that is perpendicular to the first line and passes through (-8,2)?

OpenStudy (anonymous):

Do you have any answer chioces to choose from??????

OpenStudy (anonymous):

A : Y=1/4X+4 B : Y=-4X-30 C : Y= 1/4X-2 D : Y= 4X+34